OpenStudy (anonymous):

Some help please? I've already tried it but... 1/(3x-2)+5/x=0....thanks

OpenStudy (anonymous):

Are you trying to solve for x?

OpenStudy (amistre64):

\[\frac{1}{3x-2} + \frac{5}{x}\]

OpenStudy (amistre64):

right?

OpenStudy (amistre64):

you pretty much cross multiply at this point to get x +15x -10 = 0 ---------- yadayada x +15x -10 = 0 16x -10 = 0 x = 10/16 = 5/8

OpenStudy (amistre64):

\[x(1) + 5(3x-2) = 0\]